OpenStudy (anonymous):

Limits to the infinity 2

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} {{2x^2 - 3} \over {4x^3 + 5x}}\]

OpenStudy (anonymous):

Divide the numerator and denominator by x^3, and send to infinity.

OpenStudy (anonymous):

Factor an \(x^3\) top and bottom and cancel. Then go to the limit.

OpenStudy (anonymous):

so, the answer is 0? coz the numerator > denominator?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

The numerator will go to 0, and the denominator will go to 4.

OpenStudy (anonymous):

ops, sorry invert this numerator < denominator

OpenStudy (anonymous):

no. infinity is not a number. you cannot say that infinity^2 is greater than infinite by itself

OpenStudy (anonymous):

0/4 = 0

OpenStudy (anonymous):

ah, ok!

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty}\frac{2/x-3/x^2}{4+5/x^2}=\frac{\lim_{x \rightarrow \infty}2/x+\lim_{x \rightarrow \infty}-3/x^2}{\lim_{x \rightarrow \infty}4+\lim_{x \rightarrow \infty}5/x^2}=\frac{0-0}{4+0}\]

OpenStudy (anonymous):

Right! Thanks!

OpenStudy (anonymous):

np